# Generating a Contact Matrix for Aged Care Settings in Australia: an agent-based model study

### v1

@@ 3.3 Staff and Resident Contact Analyses
Figure 4 represents the temporal distribution and intensity of contacts across agent types based on the rolling sum of contacts recorded every 10 minutes throughout the day with the contact radius of 1.5 metres. For residents, there is variability in both peaks of frequency and sustained contacts between care levels. High care residents demonstrated lower overall predicted contact frequency, with a mean rolling sum of 0.87 contacts [95% CI: 0.82 to 0.92]. The maximum recorded contact for this group was 22. Medium care residents exhibited substantial predicted contact activity throughout the day, with a mean rolling sum of 3.99 [95% CI: 3.82 to 4.16]. The maximum contact peak for this group was 66. Low care residents also showed significant predicted contact activity, with a mean rolling sum of 3.80 [95% CI: 3.63 and 3.96], with a maximum contact peak of 65. The staff contact patterns varied across the three shifts with shift 1 care staff experiencing a predicted mean rolling sum of 7.62 contacts [95% CI: 7.27 to 7.97], with a maximum contact recorded of 47. Shift 2 staff had a predicted mean rolling sum of 5.94 [95% CI: 5.64 to 6.23], with a maximum of 46. Staff in Shift 3 demonstrated the lowest predicted contact frequency, with a mean rolling sum of 0.52 [95% CI: 0.49 to 0.54], and a maximum of 15.
Figure 4: Time Series analysis of mean rolling contacts across agent types. Rolling sum of contacts between each agent type within the model calculated over 20-minute intervals with data aggregated from per-minute contact counts. Interactions are displayed for high care(a), medium care(b), and low care residents(c), as well as staff across three shifts: Shift 1(d), Shift 2(e), and Shift 3(f).
Figure 5 displays the number of predicted contacts over time by contact category, with bubble size representing the average contact duration, aggregated across 100 stochastic simulations for the 1.5m contact radius. Data were summarised in 20-minute intervals.
@@
The analysis of the activity patterns among the residents reveals several key trends, as seen in Figure 6. The data demonstrates that sitting is the most prevalent activity, occupying a significant portion of the residents’ time throughout the simulation period. This suggests that sedentary behaviour is a dominant part of daily routines, possibly indicating the need for interventions to promote more active lifestyles. Communal activities are interspersed sporadically, with noticeable simulated peaks suggesting scheduled group interactions. Predicted sleep patterns, although generally consistent, show some variability in timing, which could be attributed to personalised care schedules or differing sleep needs among residents. Predicted self-hygiene activities, while less frequent, are concentrated during specific times of the day, aligning with typical daily routines. Simulated movement between destinations appears regularly but is primarily associated with transitions between key activities such as meals and communal engagements. Interestingly, simulated mealtimes are not uniformly distributed, indicating either flexibility in meal scheduling or variations in individual preferences.
@@ 3.4 Transmission Simulation
Figure 7: Transmission network Visualisation across staff shifts. The network graphs represent simulated transmission outcomes over a 24-hour period following the introduction of a single infectious staff member per shift. Nodes correspond to individual agents, with node size proportional to the total number of contacts made by the agent. Edges represent recorded contact events, coloured red if transmission occurred and blue if transmission did not occur. Edge width reflects the uncertainty range (95% CI) of the transmission probability, and edge length is proportional to the cumulative duration of the contact between agents. Separate networks are shown for Shift 1, Shift 2, and Shift 3.
To assess the impact of booster vaccination on transmission risk, two intervention scenarios were simulated and compared against the baseline model with no vaccination. Scenario 1 implemented booster vaccination among all residents only, while Scenario 2 implemented booster vaccination among both residents and staff. Across all shifts, the introduction of vaccination reduced the number of transmission events compared to the baseline scenario, with greater reductions observed when both residents and staff were vaccinated [Figure 8].
@@
Figure 8: Transmission network visualisation under vaccination scenarios. The network graphs display simulated infection outcomes of a single infectious staff member per shift, comparing three vaccination scenarios: (i) no boosters administered (left), (ii) all residents receive a booster dose (middle), and (iii) all residents and staff receive a booster dose (right). Each row corresponds to one of three staff shifts. Nodes represent individual agents, with node size proportional to the total number of contacts recorded. Edges denote contact events between agents, with red edges indicating transmission and blue edges indicating no transmission. Edge width reflects the uncertainty range of transmission probability including vaccine effectiveness confidence intervals, and edge length is proportional to the cumulative duration of contact. Vaccination status was applied at initialisation, and transmission probability was adjusted dynamically based on scenario-specific vaccine effectiveness.
@@ 4 Discussion
This study demonstrates the value of ABMs in capturing the complex interaction patterns that traditional survey-based methods might overlook, particularly in specialised environments like aged care homes (mossong2008social; wallinga2006socialcontacts; hoang2019contactsurveys).
@@
In contrast, the simulation predicted that high care residents, who have limited mobility, would have significantly fewer interactions, primarily involving contact with staff rather than with other residents. This observation is consistent with the expected behaviour in aged care settings, where residents with higher care needs are often more isolated and dependent on staff for assistance. The reduced contact among high care residents indicates a lower likelihood of resident-to-resident transmission within this group, shifting the focus of infection control to staff-resident interactions, particularly during personal care activities. This dependency highlights the importance of ensuring that staff are adequately trained in infection prevention protocols and that there is strict adherence to hygiene practices during these interactions (mossong2008social; wallinga2006socialcontacts).
@@
Furthermore, the model’s flexibility extends to the simulation of communal spaces and their use by different resident groups at varying times. By capturing these dynamics, the model provides a more comprehensive understanding of how spatial and temporal factors influence contact patterns and potential disease transmission within a facility. This feature allows for the exploration of various "what-if" scenarios, such as the effects of altering staff shift patterns or reorganising communal spaces to minimise high-risk interactions. The insights gained from such simulations can inform the implementation of more effective infection control measures, tailored to the specific needs and configurations of each facility. Future research could build on this work by integrating the contact matrix generated in this study with a deterministic multicompartment epidemic model. This integration would enable a deeper exploration of how specific changes in contact patterns, influenced by shifts in staff schedules or the reorganisation of communal spaces, impact the transmission dynamics of infectious diseases. Such a combined approach could provide more precise estimates of the effectiveness of targeted interventions, offering a powerful tool for optimising infection control strategies in aged care settings. By linking the detailed interaction data from the ABM with the predictive capabilities of a deterministic model, researchers could advance our understanding of the interplay between behavioural dynamics and disease spread, paving the way for more informed public health decisions in these critical environments.
@@
There are several limitations to this study that warrant consideration. First is the absence of empirical validation against direct observation or sensor-based measurement of contact patterns, such as wearable RFID technology. While these technologies can offer fine-grained insight into interpersonal proximity, they present substantial challenges in aged care settings, including cost, ethical concerns, and feasibility of deployment among residents with cognitive or mobility impairments. Consequently, no such dataset currently exists for this environment. Instead, this study relied on survey of aged care workers to capture daily activity patterns, task timings, and resident interaction frequencies. Although this introduces reliance on self-reporting and inferred contact proxies, it represents a more context-specific and grounded approach than using generic population-level matrices. Future work could build on this foundation by incorporating empirical sensor data where feasible, allowing for model validation and refinement. In addition, the findings presented reflect a single run of the agent-based model under a fixed set of conditions. Given the inherent stochasticity in agent-based models, future research should consider repeated experimentation using different pseudorandom seeds to build a more comprehensive understanding of the variability in outcomes at each time point. Such an approach would allow for the assessment of the robustness of the model’s predictions and help to identify patterns that may emerge consistently across different simulations. This extension would be particularly beneficial for exploring the influence of randomness on contact dynamics and for refining the model to account for a broader range of potential scenarios.
@@ 5 Conclusion
6 Appendix A: Aged Care Survey
@@ 6.1 Study Design
A cross-sectional survey was conducted in 2023 to parameterise an agent-based model (ABM) aimed at creating a contact matrix for use in infectious disease transmission models within aged care settings in Sydney, Australia. The primary objective was to address the absence of standardised contact matrices for aged care environments, which are essential for accurately modelling disease transmission and enhancing outbreak preparedness. The survey focused on collecting comprehensive data on the interactions between staff and residents within a single aged care provider, with the goal of developing a robust contact matrix to be utilised in agent-based models.
The survey targeted aged care workers across various shifts within the facilities, gathering detailed information on their daily interactions with residents. Participants were asked to complete a structured 21-question survey designed to capture demographic data, care environment details, resident characteristics, contact and interaction patterns, and activity and mobility data. The survey’s design was informed by established methodologies that emphasise the importance of capturing both temporal aspects of daily routines and broader contextual factors in care work [crosbie2006using, nilsson2009registered, sveinsdottir2018content].
An activity diary component was included, allowing participants to log their activities and interactions in real-time, which was effective in documenting the repetitive and routine nature of care [crosbie2006using]. This provided a precise account of their daily routines. Structured questions complemented the diaries by gathering demographic data, care environment details, and resident characteristics, ensuring a comprehensive understanding of the factors influencing care practices [nilsson2009registered]. The diary entries were organised to capture relevant details about work activities, including the timing, duration, and nature of interactions with residents, aligning with recommendations to ensure thorough data collection [sveinsdottir2018content]. All responses were anonymised and securely stored, and the data were subsequently used to parameterise the agent-based model (ABM). Ethical approval for the study was obtained from the provider’s Research and Governance Office (Project code: R277) and The University of New South Wales Human Research Ethics Committee (approval number: HC230501).
@@ 6.2.1 Data Pre-parametrisation
6.3 Survey Responses and Parametrisation
6.3.1 Shifts and schedules and Staffing Levels
@@
6.3.2 Resident Care Levels and Staff Assignments
@@ 6.3.4 Task Prioritisation and Allocation
Task prioritisation, duration, and staff allocation varied according to the care level of the resident, as revealed by the survey responses. Open-ended questions such as "Can you describe your resident’s morning routine, including waking up, personal care activities, and self-care? How long do you normally spend with each resident’s self-care?" provided detailed insights into how time is allocated, where interactions take place, and which tasks are prioritised. For example, residents requiring higher levels of care often necessitate more time and assistance, especially during personal hygiene activities. The responses indicated that staff often spend significant time with these residents, sometimes requiring two staff members to assist, particularly in cases involving physical or cognitive challenges. These responses highlighted the need to allocate more staff and time to tasks involving high-care residents.
The survey responses also shed light on the environmental context of these interactions. For instance, residents with limited mobility were primarily assisted in their rooms or bathrooms, whereas those with higher mobility might engage in personal care activities in communal areas. This variation in the interaction environment highlights the importance of considering the specific locations where care is provided when allocating tasks and staff. Furthermore, structured questions regarding the activities performed during each hour of the shift provided concrete data on task durations and frequencies. For example, respondents detailed the time spent on tasks such as meal preparation, medication administration, and personal care, which can be used to estimate the average duration of these activities. Additionally, insights into staff break times and the scheduling of communal activities offer crucial data for planning staff coverage and ensuring that residents receive the necessary attention during these periods.
Finally, survey responses about the proximity and frequency of interactions between staff and residents, along with the number of residents requiring different levels of assistance during meals, help to define contact patterns. For example, staff often reported being in close proximity to residents during feeding times, with varying levels of required assistance, which can inform how close contacts are defined in terms of distance and duration.
@@ 6.3.5 Contact Definitions and Interaction Patterns
The survey provided key insights into how frequently staff interacted with residents and the proximity of these interactions, which were used to define the contact parameters within the model. Contacts were modelled based on the distance between individuals (e.g., within 1.5 or 3 meters) and the duration of the interactions, with significant contacts being defined as those lasting 3 seconds or more. This information was crucial for constructing the contact matrix within the model, which is used to simulate potential points of disease transmission within the facility. The allocation of staff to specific tasks, particularly those likely to result in close contact with residents, was also informed by the survey data. The model was parameterised to reflect which tasks, such as aiding in self-care or mealtime assistance, were most likely to involve close contact, thereby influencing the opportunities for disease transmission.
@@ 6.3.7 Environmental Setup and Population Distribution
The staff-to-resident ratios, as reported in the survey, were used to parameterise the model’s staffing levels within each care level. This ensured that the simulation reflected the workload and staff distribution, providing a realistic foundation for the model’s operations. The environmental layout and functionality within the model were also informed by the survey, particularly regarding where tasks are typically performed. The model was designed to replicate the spatial organisation of the facility, including the specific wings or areas where staff and residents are likely to interact. This spatial parametrisation was critical for accurately modelling the flow of activities and interactions within the aged care environment.

### v2

@@ 4.3 Staff and Resident Contact Analyses
To examine temporal variation in interaction intensity, we calculated a 20-minute rolling sum of contacts for each agent category. The rolling sum represents the total number of contacts involving that agent category within the preceding 20-minute window, updated at each one-minute timestep to produce a smoothed time-series of activity [Figure 4]. For residents, there is variability in both peaks of frequency and sustained contacts between care levels. High care residents demonstrated lower overall predicted contact frequency, with a mean rolling sum of 0.87 contacts [95% CI: 0.82 to 0.92]. The maximum recorded contact for this group was 22. Medium care residents exhibited substantial predicted contact activity throughout the day, with a mean rolling sum of 3.99 [95% CI: 3.82 to 4.16]. The maximum contact peak for this group was 66. Low care residents also showed significant predicted contact activity, with a mean rolling sum of 3.80 [95% CI: 3.63 and 3.96], with a maximum contact peak of 65. The staff contact patterns varied across the three shifts with shift 1 care staff experiencing a predicted mean rolling sum of 7.62 contacts [95% CI: 7.27 to 7.97], with a maximum contact recorded of 47. Shift 2 staff had a predicted mean rolling sum of 5.94 [95% CI: 5.64 to 6.23], with a maximum of 46. Staff in Shift 3 demonstrated the lowest predicted contact frequency, with a mean rolling sum of 0.52 [95% CI: 0.49 to 0.54], and a maximum of 15.
Figure 4: Rolling 20-minute contact totals for residents and staff across the day. The plots show how the number of contacts changes over time for high care (a), medium care (b), and low care residents (c), and for staff on shifts 1 (d), 2 (e), and 3 (f). Each point reflects the total number of contacts recorded in the previous 20 minutes, updated at one-minute steps. Colours indicate which agent types the contacts involved, as shown in the legend.
Predicted contact volumes varied over the course of the simulation day and differed across contactee categories. Contact counts exhibited clear temporal structure when aggregated into 20-minute intervals and averaged across stochastic simulation runs. Periods of increased contact frequency coincided with specific times of day, with variation in both the number and duration of contact events across categories [Figure 5].
@@
Across the simulated day, residents spent the majority of time in sitting or sleep states, with other activities occurring in shorter, structured intervals. Communal activities were concentrated within distinct time blocks, whereas self-hygiene events were infrequent and confined to narrow windows. Sleep activity was predominantly clustered overnight, with inter-individual variation in timing and duration. Periods of movement between destinations were brief and distributed throughout the day, typically occurring around transitions between activities. Mealtime activity was temporally dispersed across residents rather than occurring at a single synchronised time [Figure 6].
@@ 4.4 Transmission Simulation
Figure 7: Transmission networks for each staff shift following seeding of a single infectious staff member. Nodes represent individual agents and are scaled by the total number of contacts made during the simulation. Edges represent aggregated contact events between the infectious staff member and other agents, coloured red when transmission occurred and blue otherwise. Edge length corresponds to cumulative contact duration, while edge width reflects uncertainty in the estimated transmission probability (95% CI).
To assess the impact of booster vaccination on transmission risk, two intervention scenarios were simulated and compared against the baseline model, representing waned immunity with no effective protection. Scenario 1 implemented booster vaccination among all residents only, while Scenario 2 implemented booster vaccination among both residents and staff. Across all shifts, the introduction of vaccination reduced the number of transmission events compared to the baseline scenario, with greater reductions observed when both residents and staff were vaccinated [Figure 8].
@@
Figure 8: Network representation of contact and transmission outcomes during each staff shift under vaccination intervention scenarios following seeding of a single infectious staff member. Nodes correspond to individual agents and are scaled by the total number of contact events involving that agent across the simulation period. Edges represent the aggregated contact history between the infectious staff member and each contacted agent, with edge colour indicating whether transmission occurred (red) or did not occur (blue). Edge length is proportional to the cumulative duration of contact between the pair, while edge width reflects the uncertainty (95% confidence interval) associated with the estimated transmission probability after adjustment for vaccination status. Separate networks are shown for Shift 1, Shift 2, and Shift 3.
@@ 5 Discussion
This study demonstrates the value of ABMs in capturing the complex interaction patterns that traditional survey-based methods might overlook, particularly in specialised environments like aged care homes [25, 36, 17].
@@
In contrast, the simulation predicted that high care residents, who have limited mobility, would have significantly fewer interactions, primarily involving contact with staff rather than with other residents. This observation is consistent with the expected behaviour in aged care settings, where residents with higher care needs are often more isolated and dependent on staff for assistance. The reduced contact among high care residents indicates a lower likelihood of resident-to-resident transmission within this group, shifting the focus of infection control to staff-resident interactions, particularly during personal care activities. This dependency highlights the importance of ensuring that staff are adequately trained in infection prevention protocols and that there is strict adherence to hygiene practices during these interactions [25, 36].
@@
Furthermore, the model’s flexibility extends to the simulation of communal spaces and their use by different resident groups at varying times. By capturing these dynamics, the model provides a more comprehensive understanding of how spatial and temporal factors influence contact patterns and potential disease transmission within a facility. As the behavioural logic is decoupled from the specific floorplan, the framework can be applied to facilities with alternative spatial configurations, allowing these spatial dynamics to be explored in a range of architectural contexts. This feature allows for the exploration of various "what-if" scenarios, such as the effects of altering staff shift patterns or reorganising communal spaces to minimise high-risk interactions. The insights gained from such simulations can inform the implementation of more effective infection control measures, tailored to the specific needs and configurations of each facility. Future research could build on this work by integrating the contact matrix generated, in this study, with a deterministic multicompartment epidemic model. This integration would enable a deeper exploration of how specific changes in contact patterns, influenced by shifts in staff schedules or the reorganisation of communal spaces, impact the transmission dynamics of infectious diseases. Such a combined approach could provide more precise estimates of the effectiveness of targeted interventions, offering a powerful tool for optimising infection control strategies in aged care settings. By linking the detailed interaction data from the ABM with the predictive capabilities of a deterministic model, researchers could advance our understanding of the interplay between behavioural dynamics and disease spread, paving the way for more informed public health decisions in these critical environments.
@@
There are several limitations to this study that warrant consideration. First is the absence of empirical validation against direct observation or sensor-based measurement of contact patterns, such as wearable RFID technology. While these technologies can offer fine-grained insight into interpersonal proximity, they present substantial challenges in aged care settings, including cost, ethical concerns, and feasibility of deployment among residents with cognitive or mobility impairments. Consequently, no such dataset currently exists for this environment. Instead, this study relied on a survey of aged care workers to capture daily activity patterns, task timings, and resident interaction frequencies. Although this introduces reliance on self-reporting and inferred contact proxies, it represents a more context-specific and grounded approach than using generic population-level matrices. In addition, the findings presented reflect a single run of the agent-based model under a fixed set of conditions. Given the inherent stochasticity in agent-based models, repeated experimentation using different pseudorandom seeds would allow a more comprehensive understanding of variability in outcomes and support assessment of the robustness of emergent patterns. A further limitation is that the transmission component represents only primary infections within a single 24-hour period, without modelling secondary or subsequent transmission, meaning that potential multi-day outbreak trajectories are not captured in the present analysis. Finally, although Emergency Response behaviours are included in the task structure to represent acute clinical events or facility-wide disruptions, no emergency triggers were scheduled in the scenarios used here.
Future extensions of the model could incorporate scenarios with varying numbers of infectious staff across shifts, together with different staff and resident vaccination combinations. Expanding the simulation design in this way would allow systematic exploration of interactions between seeding intensity, shift structure, and immunity status, and would enable the use of regression-based analyses to quantify the relative contribution of each factor to overall transmission risk.
@@ 6 Conclusion
7 Ethics Statement
Ethics approval from The University of New South Wales Human Research Ethics Committee (approval number: HC230501) and the aged care provider’s Research and Governance Office approval (Project code: R277). Survey design, data categorisation and details on informing parameters can be found within Appendix A.
8 Code Availability
The full AnyLogic model, including the facility layout module and behavioural logic, is openly available at:
https://github.com/BREATHE-UNSW/AgedCareContactMatrix
9 Author Contribution Statement
Conceptualisation and study aims were developed by HS, DH and CRM. Survey design was performed HS, CP and MK under the supervision of DH and CRM. Ethics and the aged care research governance applications for the survey were performed by HS under the supervision of DH, CP and CRM. Model development, parameterisation, validation and calibration was performed by HS under the supervision of DH. Contact matrix generation and GLM analyses were performed by HS under the supervision of DH. Finally, this study was written by HS and reviewed by DH, MK, CP and CRM.
Appendix A provides detailed survey instruments and parameterisation rationale supporting the main Methods section.
@@ 10.1 Study Design
A cross-sectional survey was conducted in 2023 to parameterise an agent-based model (ABM) aimed at creating a contact matrix for use in infectious disease transmission models within aged care settings in Sydney, Australia. The primary objective was to address the absence of standardised contact matrices for aged care environments, which are essential for accurately modelling disease transmission and enhancing outbreak preparedness. The survey focused on collecting comprehensive data on the interactions between staff and residents within a single aged care provider, with the goal of developing a robust contact matrix to be utilised in agent-based models. The survey targeted aged care workers across various shifts within the facilities, gathering detailed information on their daily interactions with residents. Participants were asked to complete a structured 21-question survey designed to capture demographic data, care environment details, resident characteristics, contact and interaction patterns, and activity and mobility data. The survey’s design was informed by established methodologies that emphasise the importance of capturing both temporal aspects of daily routines and broader contextual factors in care work [8, 27, 33]. An activity diary component was included, allowing participants to log their activities and interactions in real-time, which was effective in documenting the repetitive and routine nature of care [8]. This provided a precise account of their daily routines. Structured questions complemented the diaries by gathering demographic data, care environment details, and resident characteristics, ensuring a comprehensive understanding of the factors influencing care practices [27]. The diary entries were organised to capture relevant details about work activities, including the timing, duration, and nature of interactions with residents, aligning with recommendations to ensure thorough data collection [33]. All responses were anonymised and securely stored, and the data were subsequently used to parameterise the agent-based model (ABM). Ethical approval for the study was obtained from the provider’s Research and Governance Office (Project code: R277) and The University of New South Wales Human Research Ethics Committee (approval number: HC230501).
@@ 10.3 Data Parameterisation
10.4 Survey Responses and Parameterisation
10.4.1 Shift Schedules and Staffing Levels
@@
10.5 Resident Care Levels and Staff Assignments
@@ 10.5.2 Task Prioritisation and Allocation
Task prioritisation, duration, and staff allocation varied according to the care level of the resident, as revealed by the survey responses. Open-ended questions such as "Can you describe your resident’s morning routine, including waking up, personal care activities, and self-care? How long do you normally spend with each resident’s self-care?" provided detailed insights into how time is allocated, where interactions take place, and which tasks are prioritised. For example, residents requiring higher levels of care often necessitate more time and assistance, especially during personal hygiene activities. The responses indicated that staff often spend significant time with these residents, sometimes requiring two staff members to assist, particularly in cases involving physical or cognitive challenges. These responses highlighted the need to allocate more staff and time to tasks involving high-care residents. The survey responses also shed light on the environmental context of these interactions. For instance, residents with limited mobility were primarily assisted in their rooms or bathrooms, whereas those with higher mobility might engage in personal care activities in communal areas. This variation in the interaction environment underscores the importance of considering the specific locations where care is provided when allocating tasks and staff. Furthermore, structured questions regarding the activities performed during each hour of the shift provided concrete data on task durations and frequencies. For example, respondents detailed the time spent on tasks such as meal preparation, medication administration, and personal care, which can be used to estimate the average duration of these activities. Additionally, insights into staff break times and the scheduling of communal activities offer crucial data for planning staff coverage and ensuring that residents receive the necessary attention during these periods. Finally, survey responses about the proximity and frequency of interactions between staff and residents, along with the number of residents requiring different levels of assistance during meals, help to define contact patterns. For example, staff often reported being in close proximity to residents during feeding times, with varying levels of required assistance, which can inform how close contacts are defined in terms of distance and duration.
@@ 10.5.3 Contact Definitions and Interaction Patterns
The survey provided key insights into how frequently staff interacted with residents and the proximity of these interactions, which were used to define the contact parameters within the model. Contacts were modelled based on the distance between individuals and the duration of the interactions, with significant contacts being defined as those lasting 3 seconds or more. This information was crucial for constructing the contact matrix within the model, which is used to simulate potential points of disease transmission within the facility. The allocation of staff to specific tasks, particularly those likely to result in close contact with residents, was also informed by the survey data. The model was parameterised to reflect which tasks, such as aiding in self-care or mealtime assistance, were most likely to involve close contact, thereby influencing the opportunities for disease transmission.
@@ 10.5.5 Environmental Setup and Population Distribution
The staff-to-resident ratios, as reported in the survey, were used to parameterise the model’s staffing levels within each care level. This ensured that the simulation reflected the workload and staff distribution, providing a realistic foundation for the model’s operations. The environmental layout and functionality within the model were also informed by the survey, particularly regarding where tasks are typically performed. The model was designed to replicate the spatial organisation of the facility, including the specific wings or areas where staff and residents are likely to interact. This spatial parameterisation was critical for accurately modelling the flow of activities and interactions within the aged care environment.